# A Cramér–Wold theorem for mixtures

### v1

@@ 1. Introduction
To apply Theorem 2.1, it is necessary to check that certain families of probability distributions are linearly independent, or equivalently, that they are identifiable. There is an extensive literature on techniques for proving identifiability. In Section 3, we review results concerning the identifiability of two particular families: the Gaussian distributions and the $t$-distributions. We also derive what we believe to be a new criterion for identifiability, based on a deep theorem in complex analysis due to Borel.
In Section 4, these ideas are applied to the important special case of Gaussian measures and Gaussian mixtures. In particular, we derive a Cramér–Wold theorem for Gaussian mixtures, Theorem 4.1, from the previously known one for Gaussian measures. It says that, if the one-dimensional projections of two mixtures of $\ell$ and $m$ normal distributions on $\mathbb{R}^{d}$ coincide on a certain pre-determined set of $(1/2)(l+m-1)(d^{2}+d-2)+1$ lines, then the two distributions are the same. A similar result, Theorem 4.2, is also established for mixtures of multivariate $t$-distributions.
We conclude in Section 5 with some remarks about potential applications of these results.
@@ 2.1. Introduction
where $\pi_{H}:\mathbb{R}^{d}\to H$ is the orthogonal projection of $\mathbb{R}^{d}$ onto $H$.
@@
  • If $\mathcal{P}$ is the family of probability measures on $\mathbb{R}^{d}$ whose supports contain at most $k$ points, and if =$\mathcal{H}=\{H_{1},\dots,H_{k+1}\}$ is any family of $k+1$ subspaces of $\mathbb{R}^{d}$ such that $H_{i}^{\perp}\cap H_{j}^{\perp}=\{0\}$ whenever $i\neq j$, then $\mathcal{H}$ is a Cramér–Wold system for $\mathcal{P}$ (Heppes, [11, Theorem $1^{\prime}$]).
@@ 2.2. Mixtures
Let $\mathcal{P}$ be a family of Borel probability measures on $\mathbb{R}^{d}$, let $\mathcal{H}$ be a collection of vector subspaces of $\mathbb{R}^{d}$, and let $\ell,m\geq 1$. Suppose that:
@@
    Each partition of $\mathcal{H}$ into $\ell+m-1$ subsets contains at least one Cramér–Wold system for $\mathcal{P}$.
Let $P$ and $Q$ be convex combinations of $\ell$ and $m$ measures from $\mathcal{P}$ respectively. If $P_{H}=Q_{H}$ for all $H\in\mathcal{H}$, then $P=Q$.
Proof.
We argue by contradiction. Suppose that $P_{H}=Q_{H}$ for all $H\in\mathcal{H}$, but that $P\neq Q$. Then the difference $P-Q$ can be written as $P-Q=\sum_{j=1}^{k}\lambda_{j}P_{j}$, where $P_{1},\dots,P_{k}\in\mathcal{P}$ are distinct, where $\lambda_{1},\dots,\lambda_{k}\in\mathbb{R}$ are non-zero, and where $k\leq\ell+m$. Set
@@
Clearly no $\mathcal{H}_{j}$ is Cramér–Wold system for $\mathcal{P}$. If we had $\cup_{j=2}^{k}\mathcal{H}_{j}=\mathcal{H}$, then we could construct a partition of $\mathcal{H}$ into $k-1$ sets (perhaps some of them empty) none of which is a Cramér–Wold system for $\mathcal{P}$, contradicting the assumption (ii) on $\mathcal{H}$. We conclude that $\cup_{j=2}^{k}\mathcal{H}_{j}\neq\mathcal{H}$, so there exists an $H_{0}\in\mathcal{H}$ such that $(P_{1})_{H_{0}}\neq(P_{j})_{H_{0}}~{}(j=2,\dots,k)$.
By assumption (i), the distinct measures in the set $\{(P_{j})_{H_{0}}:j=1,\dots d\}$ are linearly independent. In particular, since $(P_{1})_{H_{0}}\neq(P_{j})_{H_{0}}~{}(j=2,\dots,k)$, it follows that $(P_{1})_{H_{0}}$ is not in the span of $\{(P_{j})_{H_{0}}:j=2,\dots d\}$. On the other hand, we have
@@ 3.2. Linear independence of Gaussian and $t$-distributions
This result is well known. See for example [19, Example 3.1.4]. ∎
@@
Special cases of this result have previously been obtained by Otiniano et al. (case $\mu=0$), see [16, Proposition 4], and by Ho–Nguyen (case $\nu$ odd), see [12, Theorem 3.4(b)]. However, we have not been able to find the general case elsewhere in the literature, so we present a proof here.
@@
Suppose, on the contrary, that there is a linear relation of the form
(2) $\sum_{j=1}^{n}\lambda_{j}f_{\nu_{j},\mu_{j},\sigma_{j}}(x)=0\quad(x\in\mathbb{R}),$
where the triples $(\nu_{j},\mu_{j},\sigma_{j})$ are distinct and the coefficients $\lambda_{j}\neq 0$. Re-ordering the triples, we may further suppose that $\nu_{1}\geq\nu_{j}$ for all $j\geq 2$.
If all the $\nu_{j}$ are odd integers, then we may argue as follows. Each function $f_{\nu_{j},\mu_{j},\sigma_{j}}$ extends to a function in the complex plane $\mathbb{C}$
$f_{\nu_{j},\mu_{j},\sigma_{j}}(z)=c\Bigl{(}1+\frac{(z-\mu_{j})^{2}}{\nu_{j}\sigma_{j}^{2}}\Bigr{)}^{-(\nu_{j}+1)/2},$
which is holomorphic on $\mathbb{C}$ except for poles at the two points $\mu_{j}\pm i\sigma_{j}\sqrt{\nu_{j}}$. Moreover, the order of the poles at these points is exactly $(\nu_{j}+1)/2$. In particular, it follows that, if $j\geq 2$, then either $f_{\nu_{j},\mu_{j},\sigma_{j}}$ has no singularity at $\mu_{1}+i\sigma_{1}\sqrt{\nu_{1}}$, or it has a pole there of order strictly less than $\nu_{1}$. Thus
$\frac{f_{\nu_{j},\mu_{j},\sigma_{j}}(z)}{f_{\nu_{1},\mu_{1},\sigma_{1}}(z)}\to 0\quad\text{as~{}}z\to(\mu_{1}+i\sigma_{1}\sqrt{\nu_{1}})\quad(2\leq j\leq n).$
Now, by the identity principle, the relation (2) extends to
$\sum_{j=1}^{n}\lambda_{j}f_{\nu_{j},\mu_{j},\sigma_{j}}(z)=0\quad(z\in\mathbb{C},\,z\neq\mu_{j}\pm i\sigma_{j}\sqrt{\nu_{j}}).$
Dividing both sides by $f_{\nu_{1},\mu_{1},\sigma_{1}}(z)$ and then letting $z\to\mu_{1}+i\sigma_{1}\sqrt{\nu_{1}}$, we deduce that $\lambda_{1}=0$, contrary to hypothesis.
If some of the $\nu_{j}$ are even integers, then the argument needs to be adjusted slightly, since the presence of square roots means that the holomorphic extensions of $f_{\nu_{j},\mu_{j},\sigma_{j}}$ are multivalued. The problem is fixed as follows. We construct a finite set of non-intersecting half-lines terminating at the points $\mu_{j}\pm i\sigma_{j}\sqrt{\nu_{j}}$ and not crossing the real axis. The complement of these half-lines is then a simply connected domain $D$ containing $\mathbb{R}$ whose boundary contains each point $\mu_{j}\pm i\sigma_{j}\sqrt{\nu_{j}}$. Each function $f_{\nu_{j},\mu_{j},\sigma_{j}}$ has a single-valued holomorphic extension to $D$, and we can now argue exactly as before. ∎
@@ 3.3. Linear independence of measures via Borel’s theorem
Theorems 3.1 and 3.2 above are just two examples of a large class of theorems asserting the identifiability of various families of distributions. The proofs of many of these results are very similar, and indeed our proof of Theorem 3.2 relies on the same ideas. There is even an abstract formulation of the general method due to Teicher [18, Theorem 2].
In this subsection we present a further general result on identifiability/linear independence, based on a completely different principle, namely a classic result in complex analysis due to Borel [2]. We have not seen this theorem used elsewhere in the statistical literature, and we believe that the following application may be of interest.
@@
Let $\mathcal{P}$ be a family of Borel probability measures on $\mathbb{R}^{d}$. Suppose that the characteristic function of every measure in $\mathcal{P}$ is the the restriction to $\mathbb{R}^{d}$ of a nowhere-vanishing holomorphic function on $\mathbb{C}^{d}$. Then $\mathcal{P}$ is linearly independent.
@@
(iii) Suppose that characteristic function of $P$ is the restriction to $\mathbb{R}^{d}$ of a holomorphic function $f$ on $\mathbb{C}^{d}$. A sufficient condition for $f$ to be zero-free on $\mathbb{C}^{d}$ is that $P$ be infinitely divisible. This follows from [13, Theorem 3.1] and [15, Theorem 4].
(iv) The following simple example shows that the nowhere-vanishing condition in Theorem 3.3 cannot be omitted. Let $P_{0}$ and $P_{1}$ be the Dirac measures on $\mathbb{R}$ concentrated at 00 and $1$ respectively, and let $P_{2}:=(1/3)P_{0}+(2/3)P_{1}$. The characteristic functions are the $P_{j}$ are given by $\phi_{P_{0}}(\xi)=1$ and $\phi_{P_{1}}(\xi)=e^{i\xi}$ and $\phi_{P_{2}}=(1/3)(1+2e^{i\xi})$. All three are nowhere-zero on $\mathbb{R}$ and all three extend to be holomorphic on $\mathbb{C}$. However, the set $\{P_{0},P_{1},P_{2}\}$ is clearly linearly dependent. Theorem 3.3 does not apply in this case, because the holomorphic extension of $\phi_{P_{2}}$ has zeros in $\mathbb{C}$ (namely at $i\log 2+2\pi n$ for each integer $n$).
As mentioned above, the proof of Theorem 3.3 is based on a theorem of Borel. The precise version that we need is due to Green [9, p.98]:
@@ A Gaussian measure $P$ on $\mathbb{R}^{d}$ is one whose density has the form
(4) $\frac{1}{(2\pi\det(\Sigma))^{d/2}}\exp\Bigl{(}-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)\Bigr{)}\quad(x\in\mathbb{R}^{d}),$
@@ 4.2. A Cramér–Wold theorem for Gaussian mixtures
Let $S$ be a set of vectors in $\mathbb{R}^{d}$. Then the corresponding set of lines $\{\langle x\rangle:x\in S\}$ is a Cramér–Wold system for the Gaussian measures in $\mathbb{R}^{d}$ if and only if $S$ has the property that the only real symmetric $d\times d$ matrix $A$ satisfying $x^{T}Ax=0$ for all $x\in S$ is the zero matrix. A set $S$ with this property is called a symmetric-matrix uniqueness set (or sm-uniqueness set for short).
It was shown in [5] that an sm-uniqueness set for $\mathbb{R}^{d}$ spans $\mathbb{R}^{d}$ and that it contains at least $(d^{2}+d)/2$ vectors. We shall call $S$ a strong sm-uniqueness set if every subset of $S$ containing $(d^{2}+d)/2$ vectors is an sm-uniqueness set. We can now state our Cramér–Wold theorem for Gaussian mixtures.
@@
Let $P$ and $Q$ be convex combinations of $\ell$ and $m$ Gaussian measures on $\mathbb{R}^{d}$ respectively. Let $S$ be a strong sm-uniqueness set for $\mathbb{R}^{d}$ containing at least $(1/2)(\ell+m-1)(d^{2}+d-2)+1$ vectors. If $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, then $P=Q$.
@@
As for hypothesis (ii), we argue as follows. If $S$ is partitioned into $\ell+m-1$ sets, then one of them, $S_{0}$ say, must contain at least $(d^{2}+d)/2$ vectors (otherwise $S$ would contain at most $(1/2)(\ell+m-1)(d^{2}+d-2)$ vectors, contrary to assumption). As $S$ is a strong sm-uniqueness set, it follows that $S_{0}$ is an sm-uniqueness set. In summary, if $S$ is partitioned into $\ell+m-1$ sets, then at least one of them is an sm-uniqueness set. In other words, if $\mathcal{H}$ is partitioned into $\ell+m-1$ sets, then at least one of them is a Cramér–Wold system for the family of Gaussian measures. Thus hypothesis (ii) holds, and we are done. ∎
@@ 4.3. A Cramér–Wold theorem for $t$-mixtures
Let $P$ and $Q$ be convex combinations of $\ell$ and $m$ multivariate $t$-distributions on $\mathbb{R}^{d}$ respectively. Let $S$ be a strong sm-uniqueness set for $\mathbb{R}^{d}$ containing at least $(1/2)(\ell+m-1)(d^{2}+d-2)+1$ vectors. If $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, then $P=Q$.

### v2

@@ 1. Introduction
To apply Theorem 2.1, it is necessary to check that certain families of probability distributions are linearly independent, or equivalently, that they are identifiable. There is an extensive literature on techniques for proving identifiability. In Section 3, we review results concerning the identifiability of two particular families: the Gaussian distributions and the $t$-distributions. We also derive what we believe to be a new criterion for identifiability, based on a generalization of Picard’s theorem from complex analysis due to Borel.
In Section 4, these ideas are applied to the important special case of Gaussian measures and Gaussian mixtures. In particular, we derive a Cramér–Wold theorem for Gaussian mixtures, Theorem 4.1, from the previously known one for Gaussian measures. It says that, if $P$ and $Q$ are both mixtures of $m$ normal distributions on $\mathbb{R}^{d}$, and if their one-dimensional projections coincide on a certain pre-determined set of $(2m-1)(d^{2}+d-2)/2+1$ lines, then $P$ and $Q$ are the same. A similar result, Theorem 4.2, is also established for mixtures of multivariate $t$-distributions.
We conclude in Section 5 with some remarks about applications of these results.
@@ 2.1. Introduction
where $B$ is an arbitrary Borel subset of $H$ and $\pi_{H}:\mathbb{R}^{d}\to H$ is the orthogonal projection of $\mathbb{R}^{d}$ onto $H$.
@@
  • If $\mathcal{P}$ is the family of probability measures on $\mathbb{R}^{d}$ whose supports contain at most $k$ points, and if =$\mathcal{H}=\{H_{1},\dots,H_{k+1}\}$ is any family of $k+1$ subspaces of $\mathbb{R}^{d}$ such that $H_{i}^{\perp}\cap H_{j}^{\perp}=\{0\}$ whenever $i\neq j$, then $\mathcal{H}$ is a Cramér–Wold system for $\mathcal{P}$ (Heppes, [12, Theorem $1^{\prime}$]).
@@ 2.2. Mixtures
Let $\mathcal{P}$ be a family of Borel probability measures on $\mathbb{R}^{d}$, let $\mathcal{H}$ be a collection of vector subspaces of $\mathbb{R}^{d}$, and let $m\geq 1$. Suppose that:
@@
    Each partition of $\mathcal{H}$ into $12m-$ 1 subsets contains at least one Cramér–Wold system for $\mathcal{P}$.
Let $P$ and $Q$ each be convex combinations of $m$ measures from $\mathcal{P}$. If $P_{H}=Q_{H}$ for all $H\in\mathcal{H}$, then $P=Q$.
Remark.
The statement of Theorem 2.1 immediately raises the question of how to verify the hypotheses (i) and (ii). Hypothesis (i) will be treated in detail Section 3. Hypothesis (ii) will be addressed (at least in the special case of Gaussian and $t$-distributions) in Section 4.
Proof of Theorem 2.1.
We argue by contradiction. Suppose that $P_{H}=Q_{H}$ for all $H\in\mathcal{H}$, but that $P\neq Q$. Then the difference $P-Q$ can be written as $P-Q=\sum_{j=1}^{k}\lambda_{j}P_{j}$, where $P_{1},\dots,P_{k}\in\mathcal{P}$ are distinct, where $\lambda_{1},\dots,\lambda_{k}\in\mathbb{R}$ are non-zero, and where $k\leq 2m$. Set
@@
Clearly no $\mathcal{H}_{j}$ is a Cramér–Wold system for $\mathcal{P}$. If we had $\cup_{j=2}^{k}\mathcal{H}_{j}=\mathcal{H}$, then we could construct a partition of $\mathcal{H}$ into $k-1$ sets (perhaps some of them empty) none of which is a Cramér–Wold system for $\mathcal{P}$, contradicting the assumption (ii) on $\mathcal{H}$. We conclude that $\cup_{j=2}^{k}\mathcal{H}_{j}\neq\mathcal{H}$, so there exists an $H_{0}\in\mathcal{H}$ such that $(P_{1})_{H_{0}}\neq(P_{j})_{H_{0}}~{}(j=2,\dots,k)$.
By assumption (i), the distinct measures in the set $\{(P_{j})_{H_{0}}:j=1,\dots k\}$ are linearly independent. In particular, since $(P_{1})_{H_{0}}\neq(P_{j})_{H_{0}}~{}(j=2,\dots,k)$, it follows that $(P_{1})_{H_{0}}$ is not in the span of $\{(P_{j})_{H_{0}}:j=2,\dots d\}$. On the other hand, we have
@@ 3.2. Linear independence of Gaussian and $t$-distributions
This result is well known. An early reference is [20, Proposition 2]. See also [19, Example 3.1.4]. ∎
@@
See for example [13, Section 3, Example 1]. ∎
@@ 3.3. Linear independence of measures via Borel’s theorem
Theorems 3.1 and 3.2 above are just two examples of a large class of theorems asserting the identifiability of various families of distributions. The proofs of many of these results are very similar, and there is even an abstract formulation of the general method due to Teicher [18, Theorem 2].
In this subsection we present a further general result on identifiability/linear independence, based on a completely different principle, namely a generalization of Picard’s theorem from complex analysis due to Borel [2]. We have not seen this theorem used elsewhere in the statistical literature, and we believe that the following application may be of interest.
@@
Let $\mathcal{P}$ be a family of Borel probability measures on $\mathbb{R}^{d}$. Suppose that the characteristic function of every measure in $\mathcal{P}$ is the restriction to $\mathbb{R}^{d}$ of a nowhere-vanishing holomorphic function on $\mathbb{C}^{d}$. Then $\mathcal{P}$ is linearly independent.
@@
(iii) Suppose that characteristic function of $P$ is the restriction to $\mathbb{R}^{d}$ of a holomorphic function $f$ on $\mathbb{C}^{d}$. A sufficient condition for $f$ to be zero-free on $\mathbb{C}^{d}$ is that $P$ be infinitely divisible. This follows from [14, Theorem 3.1] and [16, Theorem 4].
(iv) The following simple example shows that the nowhere-vanishing condition in Theorem 3.3 cannot be omitted. Let $P_{0}$ and $P_{1}$ be the Dirac measures on $\mathbb{R}$ concentrated at 00 and $1$ respectively, and let $P_{2}:=(1/3)P_{0}+(2/3)P_{1}$. The characteristic functions are the $P_{j}$ are given by $\phi_{P_{0}}(\xi)=1$ and $\phi_{P_{1}}(\xi)=e^{i\xi}$ and $\phi_{P_{2}}=(1/3)(1+2e^{i\xi})$. All three are nowhere-zero on $\mathbb{R}$ and all three extend to be holomorphic on $\mathbb{C}$. However, the set $\{P_{0},P_{1},P_{2}\}$ is clearly linearly dependent. Theorem 3.3 does not apply in this case, because the holomorphic extension of $\phi_{P_{2}}$ has zeros in $\mathbb{C}$ (namely at $i\log 2+(2n+1)\pi$ for each integer $n$).
As mentioned above, the proof of Theorem 3.3 is based on a theorem of Borel. The precise version that we need is due to Green [10, p.98]:
@@ A Gaussian measure $P$ on $\mathbb{R}^{d}$ is one whose density has the form
(3) $\frac{1}{\det(2\pi\Sigma)^{1/2}}\exp\Bigl{(}-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)\Bigr{)}\quad(x\in\mathbb{R}^{d}),$
@@ 4.2. A Cramér–Wold theorem for Gaussian mixtures
Let $S$ be a set of vectors in $\mathbb{R}^{d}$. Then the corresponding set of lines $\{\langle x\rangle:x\in S\}$ is a Cramér–Wold system for the Gaussian measures in $\mathbb{R}^{d}$ if and only if $S$ has the property that the only real symmetric $d\times d$ matrix $A$ satisfying $x^{T}Ax=0$ for all $x\in S$ is the zero matrix. A set $S$ with this property is called a symmetric-matrix uniqueness set (or sm-uniqueness set for short). It was shown in [5] that an sm-uniqueness set for $\mathbb{R}^{d}$ spans $\mathbb{R}^{d}$ and that it contains at least $(d^{2}+d)/2$ vectors.
We shall call $S$ a strong sm-uniqueness set if every subset of $S$ containing $(d^{2}+d)/2$ vectors is an sm-uniqueness set. A method for generating examples of strong sm-uniqueness sets will be described in Section 4.4 below.
We can now state our Cramér–Wold theorem for Gaussian mixtures.
@@
Let $P$ and $Q$ each be convex combinations of $m$ Gaussian measures on $\mathbb{R}^{d}$. Let $S$ be a strong sm-uniqueness set for $\mathbb{R}^{d}$ containing at least $(2m-1)(d^{2}+d-2)/2+1$ vectors. If $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, then $P=Q$.
@@
As for hypothesis (ii), we argue as follows. If $S$ is partitioned into $12m-$ 1 sets, then one of them, $S_{0}$ say, must contain at least $(d^{2}+d)/2$ vectors (otherwise $S$ would contain at most $(2m-1)(d^{2}+d-2)/2$ vectors, contrary to assumption). As $S$ is a strong sm-uniqueness set, it follows that $S_{0}$ is an sm-uniqueness set. In summary, if $S$ is partitioned into $12m-$ 1 sets, then at least one of them is an sm-uniqueness set. In other words, if $\mathcal{H}$ is partitioned into $12m-$ 1 sets, then at least one of them is a Cramér–Wold system for the family of Gaussian measures. Thus hypothesis (ii) holds, and we are done. ∎
@@ 4.3. A Cramér–Wold theorem for $t$-mixtures
Let $P,Q$ each be convex combinations of $m$ multivariate $t$-distributions on $\mathbb{R}^{d}$. Let $S$ be a strong sm-uniqueness set for $\mathbb{R}^{d}$ containing at least $(2m-1)(d^{2}+d-2)/2+1$ vectors. If $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, then $P=Q$.
@@ 4.4. Strong sm-uniqueness sets
4.5. Two examples
In this subsection, we examine the rationale for the number $(2m-1)(d^{2}+d-2)/2+1$ appearing in Theorems 4.1 and 4.2. In particular, we explain the quadratic growth in $d$ and the linear growth in $m$.
The quadratic growth in $d$ is easily justified by the following theorem, essentially taken from [5]. We consider the special case $m=1$ (so that $P,Q$ are no longer mixtures). Thus $(2m-1)(d^{2}+d-2)/2+1=(d^{2}+d)/2$.
Theorem 4.5.
Let $d\geq 2$ and let $S$ be any set of vectors in $\mathbb{R}^{d}$ containing strictly fewer than $(d^{2}+d)/2$ vectors. Then there exist Gaussian measures $P,Q$ on $\mathbb{R}^{d}$ such that $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, but $P\neq Q$.
Proof.
By [5, Proposition 2], the fact that $S$ has fewer than $(d^{2}+d)/2$ elements implies that it is not an sm-uniqueness set. The existence of $P$ and $Q$ now follows from [5, Theorem 2]. ∎
This result explains not only the presence of the factor $(d^{2}+d)/2$ in Theorem 4.1, but also the need for considering of the notion of sm-uniqueness set. We remark in passing that the quantity $(d^{2}+d)/2$ arises here because it is the number of degrees of freedom of a $d\times d$ covariance matrix. This is discussed in more detail in [5, Section 2].
Now we turn to the dependence on $m$, namely the number of Gaussian (or $t$-) distributions making up the mixtures $P$ and $Q$. In our first example, we take $d=2$ and $m=2$. Though apparently modest, we believe that this example is instructive.
Example 4.6.
Let $S$ be the subset of $\mathbb{R}^{2}$ defined by
$S:=\Bigl{\{}{1\choose 0},{0\choose 1},{1\choose 1},{1\choose-1}\Bigr{\}}.$
Then $S$ is a strong sm-uniqueness set, and there exist measures $P,Q$ on $\mathbb{R}^{2}$, each a convex combination of two Gaussian measures, such that $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, but $P\neq Q$.
Proof.
First of all, we remark that, by [5, Corollary 3], a set of three vectors in $\mathbb{R}^{2}$ is an sm-uniqueness set provided that no two vectors in the set are multiples of one another. It follows that $S$ is a strong sm-uniqueness set.
Now set
$\Sigma_{1}:=\begin{pmatrix}2&0\\ 0&1\end{pmatrix},\quad\Sigma_{2}:=\begin{pmatrix}1&1\\ 1&2\end{pmatrix},\quad\Sigma_{1}^{\prime}:=\begin{pmatrix}2&1\\ 1&1\end{pmatrix},\quad\Sigma_{2}^{\prime}:=\begin{pmatrix}1&0\\ 0&2\end{pmatrix}.$
Let $P:=(P_{1}+P_{2})/2$, where $P_{j}$ is Gaussian with mean 00 and covariance $\Sigma_{j}$, and let $Q:=(Q_{1}+Q_{2})/2$, where $Q_{j}$ is Gaussian with mean 00 and covariance $\Sigma_{j}^{\prime}$. Clearly $P\neq Q$.
Elementary calculations show that, if $x={1\choose 0}$ or (${0\choose 1}$, then
$x^{T}\Sigma_{1}x=x^{T}\Sigma_{1}^{\prime}x\quad\text{and}\quad x^{T}\Sigma_{2}x=x^{T}\Sigma_{2}^{\prime}x,$
while if $x={1\choose 1}$ or (${1\choose-1}$, then
$x^{T}\Sigma_{1}x=x^{T}\Sigma_{2}^{\prime}x\quad\text{and}\quad x^{T}\Sigma_{2}x=x^{T}\Sigma_{1}^{\prime}x.$
Note that, in this last display, $\Sigma_{1}^{\prime}$ and $\Sigma_{2}^{\prime}$ have changed places. However, since $P$ and $Q$ are averages of the $P_{j}$ and $Q_{j}$ respectively, we do have $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$. ∎
If $d=2$ and $m=2$, then $(2m-1)(d^{2}+d-2)/2+1=7$. Example 4.6 shows that, in Theorem 4.1, this number cannot be replaced by $4$. However, calculations similar to those carried out in Example 4.6 indicate that $5$ lines do suffice. Thus, in this case at least, the bound $(2m-1)(d^{2}+d-2)/2+1$ is not sharp.
Despite the lack of sharpness, the linear growth in $m$ is correct. To see this, we consider a second example where, instead of varying the covariance matrices, we vary the means. This example is inspired by one for discrete measures presented in [12].
Example 4.7.
Let $m\geq 2$ and define a subset $S$ of $\mathbb{R}^{2}$ by
$S:=\Bigl{\{}\begin{pmatrix}\cos(j\pi/m)\\ \sin(j\pi/m)\end{pmatrix}:j=1,\dots,m\Bigr{\}}.$
Then $S$ is a strong sm-uniqueness set, and there exist measures $P,Q$ on $\mathbb{R}^{2}$, each a convex combination of $m$ Gaussian measures, such that $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, but $P\neq Q$.
Proof.
As in the previous example, a set of three vectors in $\mathbb{R}^{2}$ is an sm-uniqueness set provided that no two vectors in the set are multiples of one another. It follows that $S$ is a strong sm-uniqueness set.
For $j=1,\dots,m$, set
$\mu_{j}:=\begin{pmatrix}\cos((2j-1/2)\pi/m)\\ \sin((2j-1/2)\pi/m)\end{pmatrix}\quad\text{and}\quad\mu_{j}^{\prime}:=\begin{pmatrix}\cos((2j+1/2)\pi/m)\\ \sin((2j+1/2)\pi/m)\end{pmatrix}.$
Thus the sequence $\mu_{1},\mu_{1}^{\prime},\mu_{2},\mu_{2}^{\prime},\dots,\mu_{m},\mu_{m}^{\prime}$ traces out, in order, the vertices of a regular $(2m)$-gon centred at the origin. The lines defined by the set $S$ are precisely the bisectors of this $(2m)$-gon passing through the midpoints of the edges. For $j=1,\dots,m$, let $P_{j},Q_{j}$ be Gaussian measures with means at $\mu_{j},\mu_{j}^{\prime}$ respectively, all of them having covariance matrix equal to the identity. Finally, let $P=(1/m)\sum_{j=1}^{m}P_{j}$ and $Q:=(1/m)\sum_{j=1}^{m}Q_{j}$. Then it is geometrically clear that $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, but evidently $P\neq Q$. ∎
This example works because of the symmetry of the set-up. Though such symmetry is rather exceptional, there is nothing in Theorem 4.1 to exclude this possibility, and so it needs to be taken into account when deciding the number of lines to be used. Example 4.7 shows that (when $d=2$), this is number has to be at least $m+1$. This explains the linear growth in $m$ in the bound $(2m-1)(d^{2}+d-2)/2+1$.